OpenStudy (swaugh):

medal + fan What is the length of line segment A B if A is point one comma one and B is point nine comma seven? The point A is plotted at one comma one and point B is plotted at nine comma seven on a coordinate plane. The points are connected with a diagonal line segment. 8 ten twelve fourteen

alones (alones):

hey do u have 15 there ?? or is that all

OpenStudy (swaugh):

thats all

jimthompson5910 (jim_thompson5910):

Use the distance formula \[\Large d = \sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]

jimthompson5910 (jim_thompson5910):

\[\Large A = (x_1,y_1) = (1,1)\] \[\Large B = (x_2,y_2) = (9,7)\]

OpenStudy (swaugh):

i dont understand how to use that formula

OpenStudy (swaugh):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Plug in the given coordinates and simplify \[\Large d = \sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] \[\Large d = \sqrt{(9-1)^2+(7-1)^2}\] \[\Large d = \sqrt{8^2+6^2}\] \[\Large d = ???\]

OpenStudy (swaugh):

i got 44 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

8^2 is equal to what?

jimthompson5910 (jim_thompson5910):

8^2 means 8*8 btw 6^2 = 6*6

OpenStudy (swaugh):

8^2 = 64

OpenStudy (swaugh):

6^2= 36

jimthompson5910 (jim_thompson5910):

so we then will have 64+36 = 100 now take the square root of 100 to get ??

OpenStudy (swaugh):

10 is my answer righttt

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (swaugh):

thank you so much

jimthompson5910 (jim_thompson5910):

np